# disposition

### v2

@@ E.3 Additional Lemmas
For the test for the sign of $\sigma$ involved in the construction of confidence intervals for ${C}$ as explained in Subsection 6.2 and in the proof of Proposition 6.5, we need a central limit theorem for the empirical covariance $\widehat{\sigma}_{n}$, which we present here.
@@
For $\mathsf{r}_{1}$ and $\mathsf{r}_{2}$, we get complex solutions with a non-zero imaginary part for any $\mathsf{q}(\mathsf{q}-1)+\mathsf{p}(\mathsf{p}-1)+\frac{1}{4}<0$, which is equivalent to $(\mathsf{p},\mathsf{q})\in U:=\{x\in\mathbb{R}^{2}:\|x-(0.5,0.5)\|_{2}<0.5\}$. Here, $U$ is the open disk centered around $(0.5,0.5)$ with a radius of $0.5$, shown in red in Figure 10. We do not have to consider these complex solutions as they are not in $\mathbb{D}$. We further define the closure of the open disk $U$ as $\overline{U}$, and its border as $\partial U:=\overline{U}\backslash U$.
For $(\mathsf{p},\mathsf{q})\in[0,1]^{2}\backslash U$, we distinguish the following four cases that correspond to the four colored areas in Figure 10. In each case, we show that the possible roots $\mathsf{r}_{1}$ and $\mathsf{r}_{2}$ are not in $\mathbb{D}$.
$\bullet$ Case 1: Let $(\mathsf{p},\mathsf{q})\in[0,0.5]^{2}\backslash U$ (dark blue area in Figure 10). For $\mathsf{p}=0$, i.e., for any point on the left edge of the blue area, we get
@@
Furthermore, for any $(\mathsf{p},\mathsf{q})\in[0,0.5]^{2}\backslash\overline{U}$, it holds that
@@
since $\mathsf{p}+\sqrt{\mathsf{q}(\mathsf{q}-1)+\mathsf{p}(\mathsf{p}-1)+\frac{1}{4}}\leq\mathsf{p}+\sqrt{\mathsf{p}(\mathsf{p}-1)+\frac{1}{4}}=\mathsf{p}+|\mathsf{p}-\frac{1}{2}|=\mathsf{p}-\mathsf{p}+\frac{1}{2}=\frac{1}{2}$. Hence, when moving from the left edge where $\mathsf{p}=0$ towards the right in the blue area, $\mathsf{r}_{1}$ decreases (weakly) from a starting value of zero such that we can conclude that $\mathsf{r}_{1}\leq 0$ for all $(\mathsf{p},\mathsf{q})\in[0,0.5]^{2}\backslash\overline{U}$.
@@
Hence, for any $(\mathsf{p},\mathsf{q})\in[0,0.5]^{2}\backslash U$, we conclude that $\mathsf{r}_{1}\leq 0=\max(0,\mathsf{p}+\mathsf{q}-1)$. As $\mathsf{r}_{2}\leq\mathsf{r}_{1}$ (for any real-valued solution in $\mathbb{D}$) the same claim follows directly for $\mathsf{r}_{2}$.
$\bullet$ Case 2: Let $(\mathsf{p},\mathsf{q})\in\big{(}[0,0.5]\times[0.5,1]\big{)}\backslash U$ (orange area in Figure 10). For $\mathsf{p}=0$, we get
@@
Furthermore, for any $(\mathsf{p},\mathsf{q})\in\big{(}[0,0.5]\times[0.5,1]\big{)}\backslash\overline{U}$, it holds that
@@
and $\mathsf{p}+\sqrt{\mathsf{q}(\mathsf{q}-1)+\mathsf{p}(\mathsf{p}-1)+\frac{1}{4}}\leq\mathsf{p}+\sqrt{\mathsf{p}(\mathsf{p}-1)+\frac{1}{4}}=\mathsf{p}+|\mathsf{p}-\frac{1}{2}|=\mathsf{p}-\mathsf{p}+\frac{1}{2}=\frac{1}{2}$. Hence, when moving from the left edge ($\mathsf{p}=0$) towards the right in the orange area, $\mathsf{r}_{2}$ grows faster than the upper bound $\mathsf{p}=\min(\mathsf{p},\mathsf{q})$ in $\mathbb{D}$, such that we can conclude that $\mathsf{r}_{2}\geq\min(\mathsf{p},\mathsf{q})$ for all $(\mathsf{p},\mathsf{q})\in\big{(}[0,0.5]\times[0.5,1]\big{)}\backslash\overline{U}$.
@@
Hence, for any $(\mathsf{p},\mathsf{q})\in\big{(}[0,0.5]\times[0.5,1]\big{)}\backslash U$, we conclude that $\mathsf{r}_{2}\geq\min(\mathsf{p},\mathsf{q})$. As $\mathsf{r}_{1}\geq\mathsf{r}_{2}$ (for any real-valued solutions in $\mathbb{D}$) the same claim holds for $\mathsf{r}_{1}$.
$\bullet$ Case 3: Let $(\mathsf{p},\mathsf{q})\in\big{(}[0.5,1]\times[0,0.5]\big{)}\backslash U$ (yellow area in Figure 10). The argument in this case is equivalent to Case 2 by simply interchanging $\mathsf{p}$ and $\mathsf{q}$.
$\bullet$ Case 4: $(\mathsf{p},\mathsf{q})\in[0.5,1]^{2}\backslash U$ (cyan area in Figure 10). For the right edge where $\mathsf{p}=1$, we get
@@
Furthermore, for $(\mathsf{p},\mathsf{q})\in[0.5,1]^{2}\backslash\overline{U}$, it holds that
@@
and $\mathsf{p}-\sqrt{\mathsf{q}(\mathsf{q}-1)+\mathsf{p}(\mathsf{p}-1)+\frac{1}{4}}\geq\mathsf{p}-\sqrt{\mathsf{p}(p-1)+\frac{1}{4}}=\mathsf{p}-|\mathsf{p}-\frac{1}{2}|=\mathsf{p}-\mathsf{p}+\frac{1}{2}=\frac{1}{2}$. Hence, when moving from the right edge ($\mathsf{p}=1$) towards the left in the cyan area, $\mathsf{r}_{1}$ decreases faster than the lower bound $\max(0,\mathsf{p}+\mathsf{q}-1)$ in $\mathbb{D}$, such that we can conclude that $\mathsf{r}_{1}\leq\max(0,\mathsf{p}+\mathsf{q}-1)$ for all $(\mathsf{p},\mathsf{q})\in[0.5,1]^{2}\backslash\overline{U}$.
@@ Appendix F Additional Tables and Graphs
Figure 11: This figure depicts Cole’s ${C}$ with a fixed value of 0.3 and the corresponding values of Yule’s ${Q}$ and the phi coefficient for all combinations of marginal event probabilities $\mathbb{P}(A)$ and $\mathbb{P}(B)$.
Figure 12: This figure depicts Cole’s ${C}$ with a fixed value of -0.7 and the corresponding values of Yule’s ${Q}$ and the phi coefficient for all combinations of marginal event probabilities $\mathbb{P}(A)$ and $\mathbb{P}(B)$.
@@
alcohol (ALC) $12{,}922$
cigarettes (CIG) $13{,}176$
amphetamines (AMP) $13{,}147$
tranquilizers (TRQ) $13{,}028$
narcotics (NAR) $12{,}883$
LSD (LSD) $10{,}980$
psychedelics (PSY) $10{,}975$
sedatives (SED) $13{,}088$
MDMA (MDM)   $\;\,6{,}552$
cocaine (COK) $13{,}024$
crack (CRA) $12{,}505$
methamphetamines (MET)   $\;\,4{,}313$
heroin (HER) $13{,}005$
@@
alcohol (ALC) $4{,}220$
marijuana (MAR) $4{,}313$
cigarettes (CIG) $4{,}372$
amphetamines (AMP) $4{,}354$
tranquilizers (TRQ) $4{,}414$
narcotics (NAR) $4{,}423$
LSD (LSD) $2{,}163$
psychedelics (PSY) $2{,}170$
sedatives (SED) $4{,}413$
MDMA (MDM) $4{,}410$
cocaine (COK) $4{,}408$
crack (CRA) $4{,}428$
heroin (HER) $4{,}421$

### v3

@@ E.3 Additional Lemmas
For the test for the sign of $\sigma$ involved in the construction of confidence intervals for C{C} as explained in Subsection 6.2 and in the proof of Proposition 6.5, we need a central limit theorem for the empirical covariance $\widehat{\sigma}_{n}$, which we present here.
@@
For $\mathsf{r}_{1}$ and $\mathsf{r}_{2}$, we get complex solutions with a non-zero imaginary part for any $\mathsf{q}(\mathsf{q}-1)+\mathsf{p}(\mathsf{p}-1)+\frac{1}{4}<0$, which is equivalent to $(\mathsf{p},\mathsf{q})\in U:=\{x\in\mathbb{R}^{2}:\|x-(0.5,0.5)\|_{2}<0.5\}$. Here, $U$ is the open disk centered around $(0.5,0.5)$ with a radius of $0.5$, shown in red in Figure 10. We do not have to consider these complex solutions as they are not in $\mathbb{D}$. We further define the closure of the open disk $U$ as $\overline{U}$, and its border as $\U\partial U:=\overline{U}\backslash$ U.
For $\U(\mathsf{p},\mathsf{q})\in[0,1]^{2}\backslash$ U, we distinguish the following four cases that correspond to the four colored areas in Figure 10. In each case, we show that the possible roots $\mathsf{r}_{1}$ and $\mathsf{r}_{2}$ are not in $\mathbb{D}$.
$\bullet$ Case 1: Let $\U(\mathsf{p},\mathsf{q})\in[0,0.5]^{2}\backslash$ U (dark blue area in Figure 10). For $\mathsf{p}=0$, i.e., for any point on the left edge of the blue area, we get
@@
Furthermore, for any $\U¯(\mathsf{p},\mathsf{q})\in[0,0.5]^{2}\backslash\overline${U}, it holds that
@@
since $\mathsf{p}+\sqrt{\mathsf{q}(\mathsf{q}-1)+\mathsf{p}(\mathsf{p}-1)+\frac{1}{4}}\leq\mathsf{p}+\sqrt{\mathsf{p}(\mathsf{p}-1)+\frac{1}{4}}=\mathsf{p}+|\mathsf{p}-\frac{1}{2}|=\mathsf{p}-\mathsf{p}+\frac{1}{2}=\frac{1}{2}$. Hence, when moving from the left edge where $\mathsf{p}=0$ towards the right in the blue area, $\mathsf{r}_{1}$ decreases (weakly) from a starting value of zero such that we can conclude that $\mathsf{r}_{1}\leq 0$ for all $\U¯(\mathsf{p},\mathsf{q})\in[0,0.5]^{2}\backslash\overline${U}.
@@
Hence, for any $\U(\mathsf{p},\mathsf{q})\in[0,0.5]^{2}\backslash$ U, we conclude that $\mathsf{r}_{1}\leq 0=\max(0,\mathsf{p}+\mathsf{q}-1)$. As $\mathsf{r}_{2}\leq\mathsf{r}_{1}$ (for any real-valued solution in $\mathbb{D}$) the same claim follows directly for $\mathsf{r}_{2}$.
$\bullet$ Case 2: Let $\U(\mathsf{p},\mathsf{q})\in\big([0,0.5]\times[0.5,1]\big)\backslash$ U (orange area in Figure 10). For $\mathsf{p}=0$, we get
@@
Furthermore, for any $\U¯(\mathsf{p},\mathsf{q})\in\big([0,0.5]\times[0.5,1]\big)\backslash\overline${U}, it holds that
@@
and $\mathsf{p}+\sqrt{\mathsf{q}(\mathsf{q}-1)+\mathsf{p}(\mathsf{p}-1)+\frac{1}{4}}\leq\mathsf{p}+\sqrt{\mathsf{p}(\mathsf{p}-1)+\frac{1}{4}}=\mathsf{p}+|\mathsf{p}-\frac{1}{2}|=\mathsf{p}-\mathsf{p}+\frac{1}{2}=\frac{1}{2}$. Hence, when moving from the left edge ($\mathsf{p}=0$) towards the right in the orange area, $\mathsf{r}_{2}$ grows faster than the upper bound $\mathsf{p}=\min(\mathsf{p},\mathsf{q})$ in $\mathbb{D}$, such that we can conclude that $\mathsf{r}_{2}\geq\min(\mathsf{p},\mathsf{q})$ for all $\U¯(\mathsf{p},\mathsf{q})\in\big([0,0.5]\times[0.5,1]\big)\backslash\overline${U}.
@@
Hence, for any $\U(\mathsf{p},\mathsf{q})\in\big([0,0.5]\times[0.5,1]\big)\backslash$ U, we conclude that $\mathsf{r}_{2}\geq\min(\mathsf{p},\mathsf{q})$. As $\mathsf{r}_{1}\geq\mathsf{r}_{2}$ (for any real-valued solutions in $\mathbb{D}$) the same claim holds for $\mathsf{r}_{1}$.
$\bullet$ Case 3: Let $\U(\mathsf{p},\mathsf{q})\in\big([0.5,1]\times[0,0.5]\big)\backslash$ U (yellow area in Figure 10). The argument in this case is equivalent to Case 2 by simply interchanging $\mathsf{p}$ and $\mathsf{q}$.
$\bullet$ Case 4: $\U(\mathsf{p},\mathsf{q})\in[0.5,1]^{2}\backslash$ U (cyan area in Figure 10). For the right edge where $\mathsf{p}=1$, we get
@@
Furthermore, for $\U¯(\mathsf{p},\mathsf{q})\in[0.5,1]^{2}\backslash\overline${U}, it holds that
@@
and $\mathsf{p}-\sqrt{\mathsf{q}(\mathsf{q}-1)+\mathsf{p}(\mathsf{p}-1)+\frac{1}{4}}\geq\mathsf{p}-\sqrt{\mathsf{p}(p-1)+\frac{1}{4}}=\mathsf{p}-|\mathsf{p}-\frac{1}{2}|=\mathsf{p}-\mathsf{p}+\frac{1}{2}=\frac{1}{2}$. Hence, when moving from the right edge ($\mathsf{p}=1$) towards the left in the cyan area, $\mathsf{r}_{1}$ decreases faster than the lower bound $\max(0,\mathsf{p}+\mathsf{q}-1)$ in $\mathbb{D}$, such that we can conclude that $\mathsf{r}_{1}\leq\max(0,\mathsf{p}+\mathsf{q}-1)$ for all $\U¯(\mathsf{p},\mathsf{q})\in[0.5,1]^{2}\backslash\overline${U}.
@@ Appendix F Additional Tables and Graphs
Figure 11: This figure depicts Cole’s C{C} with a fixed value of 0.3 and the corresponding values of Yule’s Q{Q} and the phi coefficient for all combinations of marginal event probabilities $\mathbb{P}(A)$ and $\mathbb{P}(B)$.
Figure 12: This figure depicts Cole’s C{C} with a fixed value of -0.7 and the corresponding values of Yule’s Q{Q} and the phi coefficient for all combinations of marginal event probabilities $\mathbb{P}(A)$ and $\mathbb{P}(B)$.
@@
alcohol (ALC) 12,92212{,}922
cigarettes (CIG) 13,17613{,}176
amphetamines (AMP) 13,14713{,}147
tranquilizers (TRQ) 13,02813{,}028
narcotics (NAR) 12,88312{,}883
LSD (LSD) 10,98010{,}980
psychedelics (PSY) 10,97510{,}975
sedatives (SED) 13,08813{,}088
MDMA (MDM)   6,552\;\,6{,}552
cocaine (COK) 13,02413{,}024
crack (CRA) 12,50512{,}505
methamphetamines (MET)   4,313\;\,4{,}313
heroin (HER) 13,00513{,}005
@@
alcohol (ALC) 4,2204{,}220
marijuana (MAR) 4,3134{,}313
cigarettes (CIG) 4,3724{,}372
amphetamines (AMP) 4,3544{,}354
tranquilizers (TRQ) 4,4144{,}414
narcotics (NAR) 4,4234{,}423
LSD (LSD) 2,1632{,}163
psychedelics (PSY) 2,1702{,}170
sedatives (SED) 4,4134{,}413
MDMA (MDM) 4,4104{,}410
cocaine (COK) 4,4084{,}408
crack (CRA) 4,4284{,}428
heroin (HER) 4,4214{,}421